If the magnitude of the sum of two vectors is less than the magnitude of either vector, then:


1
Expert's answer
2020-11-16T06:40:46-0500

The magnitude of the sum of two vectors can be found from the cosine law applied to "\triangle AOB":


"c = \sqrt{a^2 + b^2 - 2ab\cos(180-\theta)} = \sqrt{a^2 + b^2 + 2ab\cos\theta}"

where "\theta" is the angle between the vectors "\mathbf{a}" and "\mathbf{b}".

Thus, it should be "\theta >90\degree" for "c<a" and "c<b".


Answer. If the magnitude of the sum of two vectors is less than the magnitude of either vector, then the angle between these two vectors is obtuse (">90\degree" ).